# Phase space analysis of CCDM cosmologies

### v1

@@ III.4 Model: $\Gamma=\Gamma_{0}H^{-1}$
In this series of matter creation models, one of the interesting matter creation rate is $\Gamma=\Gamma_{0}H^{-1}$, where $\Gamma_{0}$ is constant having dimension equal to the dimension of the square of the Hubble rate. In the following we describe the influence of this model on both one-fluid and two-fluid systems.
III.4.1 One fluid system
The dynamics of the matter creation model in the one-fluid system is described by the following differential equation
$z^{\prime}=\frac{3}{2}z\left(1-z-\frac{\gamma z^{2}}{3(1-z)}\right)$ (35)
where $\gamma=\Gamma_{0}/H_{0}^{2}$ is a positive dimensionless constant and as one can notice the above equation admits one singularity at $z=1$. Therefore, we regularize the above equation and after regularization, it becomes
$z^{\prime}=\frac{3}{2}z\left((1-z)^{2}-\frac{\gamma z^{2}}{3}\right).$ (36)
The decelerating parameter $q=\frac{1}{2}\left[1-\frac{\gamma z^{2}}{(1-z)^{2}}\right]$ at present epoch (i.e. $z=1/2)$ takes the form $q=\frac{1-\gamma}{2}$ from which one derives that at present time accelerating expansion of the universe needs $\gamma>1$. We evaluate the critical points of this one dimensional system and in Table 4 we present them along with their qualitative features. Each critical point represents an ordered pair where the first component denotes the DM density parameter and the second component is related to the $z$ variable. The critical point $D^{\prime}_{0}(1,0)$ is always unstable and it represents matter dominated decelerating phase. Again, $D^{\prime}_{1}\left(1,\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$ represents the matter dominated accelerating attractor. The one dimensional vector field involved in the equation (36) is displayed in the upper left plot of Fig. 7. This model gives the transition of our universe from its past $(H\longrightarrow+\infty)$ decelerating phase to the current (or at finite time where $H$ adopts finite value) accelerating phase.
III.4.2 Two fluid system
@@
where $\gamma=\Gamma_{0}/H_{0}^{2}$ is again a dimensionless parameter and it is positive. As the system admits a singularity at $z=1$, hence, we regularize this system and after its regularization, it reduces to
$x^{\prime}=$ $x(1-x)\left(3w(1-z)^{2}+\gamma z^{2}\right),$ (38a)
$z^{\prime}=$ $\frac{3}{2}z(1-z)\left[\left(1+w(1-x)\right)(1-z)^{2}-\frac{\gamma xz^{2}}{3}\right].$ (38b)
The Friedmann constraint and the equation (24) ensure that the physical space is unit square, namely, $R_{4}=[0,1]^{2}$. If we look into the system (38), we find that $x=0,~{}x=1,~{}z=0,~{}z=1$ are invariant manifolds. So, the domain $R_{4}$ is positively invariant. For this model we obtain the decelerating parameter as $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\gamma xz^{2}}{(1-z)^{2}}\right]$. Thus, at present time (i.e. $z=1/2$) to obtain an accelerating phase of the universe the viable orbits should maintain the relation $(1-\gamma)x+(1+3w)(1-x)<0$. The critical points and different cosmological parameter at these points are depicted in the Table 4. Now we shall analyze the qualitative features of the system (38) for different values of $w$.
One fluid system
Critical point Existence Stability $\Omega_{\rm dm}$ $q$ Acceleration
$D^{\prime}_{0}(1,0)$ Always Always Unstable 1 $\frac{1}{2}$ No
$D^{\prime}_{1}\left(1,\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$ $\gamma>0$ Always Stable 1 $-1$ Yes
Two fluid system
Critical point Existence Eigenvalue Stability $\Omega_{f}$ $\Omega_{\rm dm}$ $q$ Acceleration
$D_{0}(0,0)$ Always $\left(3w,\frac{3}{2}(1+w)\right)$ Stable if $w<-1$; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if $-1<w<0$;
@@
$D_{2}(0,1)$ Always $(\gamma,0)$ Non-hyperbolic Saddle if $w>-1$; 1 0 Undefined Undetermined
Unstable if $w<-1$
$D_{3}(1,1)$ Always $\left(-\gamma,\frac{\gamma}{2}\right)$ Always Saddle 0 1 $\infty$- ∞ Yes
$D_{4}\left(1,\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$ $\gamma>0$ $\left(-\frac{\gamma(1+w)}{\left(1+\sqrt{\frac{\gamma}{3}}\right)^{2}},-\frac{\gamma}{\left(1+\sqrt{\frac{\gamma}{3}}\right)^{2}}\right)$ Stable if $w>-1$ 0 1 $-1$ Yes
Saddle if $w<-1$
$D_{5}\left(0,z_{c}\right)$ $w=-1$ $\left(\gamma z_{c}^{2}-3(1-z_{c})^{2},0\right)$ Stable if $z_{c}<\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ $1$ 00 $-1$ Yes
@@
$D_{6}\left(x_{c},\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$ $w=-1$ $\left(0,-\frac{\gamma x_{c}}{\left(1+\sqrt{\frac{\gamma}{3}}\right)^{2}}\right)$ Stable $1-x_{c}$ $x_{c}$ $-1$ Yes
@@
Table 4: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (36) and (38) with the matter creation rate $\Gamma=\Gamma_{0}H^{-1}$.
@@
Figure 7: Description of the phase space controlled by the matter creation rate $\Gamma=\Gamma_{0}H^{-1}$, for both the one and two-fluid systems. Upper Left Plot: This is the vector plot of the one dimensional system (36) considering $\gamma=3$. However, for other positive values of $\gamma$, we can find similar type of plot. Upper Right Plot: The phase plot of the system (38) when we have assumed $w=1.5$ and $\gamma=2$. For other values of $w>0$ and $\gamma>0$, we can also obtain similar phase space structure. Lower Left Plot: The phase space of the system (38) when the equation of state $w$ takes the value 00. Here we use $\gamma=2$ but any positive value of $\gamma$ gives same type of phase portrait. Lower Right Plot: The phase plot of the system (38) considering $w=-0.8$ and $\gamma=3$. Also, we can get similar type of graphics for any positive value of $\gamma$ and negative value of $w$ in the interval $(-1,0)$. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). The black dotted curve separating the cyan and yellow regions corresponds to $q=-1$.
Figure 8: This figure also depicts the phase space of the two-fluid system driven by the matter creation rate $\Gamma=\Gamma_{0}H^{-1}$ but here we consider the case when $w\leq-1$. Left Plot: The phase space of the system (38) when the equation of state $w$ takes the value $-1$. Here we use $\gamma=2$ but any positive value of $\gamma$ gives similar type of phase portrait. Right Plot: The phase plot of the system (38) when we assume $w=-1.5$ and $\gamma=3$. For other values of $w<-1$ and $\gamma>0$, we can also obtain similar graphics. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). The black dotted curve (in the left graph this curve is not visible because of the presence of the critical line) separating the cyan and yellow regions corresponds to $q=-1$.
  1. 1.
@@
    When $w>0$: When the equation of state of the second fluid adopts positive values, we obtain five isolated critical points, namely, $D_{0},~{}D_{1},~{}D_{2},~{}D_{3}$ and $D_{4}$ in which $D_{0},~{}D_{2}$ are dominated by the second fluid only ($\Omega_{f}=1$) and $D_{1},~{}D_{3},~{}D_{4}$ are DM dominated points ($\Omega_{\rm dm}=1$). Inspecting the nature of eigenvalues from Table 4, according to the linear stability analysis, we find that $D_{0}$ is unstable, $D_{1},~{}D_{3}$ both are saddle by nature and $D_{4}$ is the only globally stable point. The critical point $D_{2}$ at which decelerating parameter is undefined, behaves like saddle point because along $z$ axis, $z$ is increasing and along $z=1$ line $x$ is increasing. Here, the points $D_{0},~{}D_{1}$ lie in the decelerating phase and $D_{3},D_{4}$ also belong to the accelerated phase of the universe. The phase space trajectories are highlighted in the upper right plot of the Fig. 7, indicating a transition of our universe from the second fluid dominated decelerating phase to a DM dominated accelerating phase where the slowing down of the cosmic acceleration is observed similar to what we have already seen for the two-fluid system of the Model $\Gamma=\Gamma_{0}$ with $w>0$, see section III.1.
  2. 2.
@@
    When $w=0$: In this case, we obtain total three isolated critical points $D_{2},~{}D_{3},~{}D_{4}$ and one critical line $D_{7}(x_{c},0)$. In our phase space, $x^{\prime}$ is always positive and, $z$ is increasing for $z<\frac{1}{1+\sqrt{\frac{\gamma}{3}}\sqrt{x}}$ and it is decreasing for $z>\frac{1}{1+\sqrt{\frac{\gamma}{3}}\sqrt{x}}$. Therefore, $D_{7}$ where DM and DE coexist, is always unstable and $D_{2},~{}D_{3}$ give saddle type behavior. Since in our domain $D_{4}$ is the only stable point, we claim that it is a globally stable point. The accelerated phase includes $D_{3},~{}D_{4}$ and on the other side, the decelerated phase comprises the line $D_{7}$. The phase space orbits are depicted in the lower left plot of the Fig. 7, showing an alteration of the universe from the decelerating scaling solutions (DM and second fluid both exist in the picture) to DM dominated accelerating era where the slowing down of the cosmic acceleration is noted -- the same feature has been observed in the model $\Gamma=\Gamma_{0}$ with $w=0$, see section III.1.
  3. 3.
@@
    When $-\frac{1}{3}\leq w<0$: In this interval of $w$, once again we can get five isolated critical points, namely, $D_{0},~{}D_{1},~{}D_{2},~{}D_{3}$ and $D_{4}$ by solving the autonomous system (38). Looking into the eigenvalues from the Table 3 and using the direction of the flow on the boundary of phase space, one can easily conclude about the stability of the critical points. Therefore, $D_{0},~{}D_{2}$ and $D_{3}$ give saddle type character, $D_{1}$ shows unstable nature and as before $D_{4}$ corresponds to matter dominated late time globally stable point. Here also, $D_{0}$, $D_{1}$ lie in the decelerating phase and $D_{3}$ belongs to the accelerating phase. The evolution of the solution curves are clearly observed in the lower right plot of the Fig. 7, highlighting the alteration from DM dominated decelerating epoch to DM dominated accelerating epoch with its slowing down nature, which one can meanwhile find in the Model III.1 with the case $-1/3<w<0$.
  4. 4.
    When $-1<w<-\frac{1}{3}$: This case properly replicates the above case where range of $w$ is $-\frac{1}{3}\leq w<0$. The only difference lie on the decelerating parameter at the point $D_{0}$ which shows acceleration. Thus the phase plot is the lower right plot of the Fig. 7.
  5. 5.
@@
    When $w=-1$: Here, looking into the dynamical system (38), we can conclude that we have a set of critical points which contains two isolated critical points $D_{1},~{}D_{3}$ and two critical lines $D_{5}(0,z_{c})$, $D_{6}\left(x_{c},\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$. In the phase space above the line $z=\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$, $x$ and $z$ show increasing and decreasing character, respectively, and below this line, $x$ is decreasing function while $z$ is increasing function. The equation of the separatrix joining the points $D_{1}$ and $\left(0,\frac{1}{\sqrt{\frac{\gamma}{3}}}\right)$ is $z=\sqrt{\frac{3}{\gamma}}(1-z)\sqrt{1-x}$ which separates the phase space below the line $z=\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ into two parts; one is below the separatrix and other one is the above the separatrix. In the first part, any trajectory approaches to the line $D_{5}$ and in the second part, trajectories converge to the line $D_{6}$. Therefore, $A_{6}$ where DM and DE coexist (except the end points), gives stable like nature. Here, $D_{1}$ and $D_{6}$ with $z>\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ show unstable nature and $D_{3}$ is a saddle point. The accelerating phase include $D_{3}$, $D_{5}$, $D_{6}$ and $D_{1}$ corresponds to deceleration. The part of the line $D_{5}$ below $z=\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ correspond to late time completely DE dominated stable points. On the other hand, the line $D_{6}$ has the ability to solve the coincidence problem. The phase space structure is properly presented in the left plot of the Fig. 8 which depicts two types of transitions: the first one is DM dominated decelerating phase to accelerating scaling solutions (DM and DE both exist) and the second one is DM dominated decelerating phase to completely DE dominated accelerating phase, that is similar to the case $w=-1$ of the Model III.1.
  6. 6.
@@
    When $w<-1$: In the phantom region of $w$, we have again five critical points, namely, $D_{0},~{}D_{1},~{}D_{2},~{}D_{3},~{}D_{4}$ which lie on the boundary of the phase space region $R_{4}$. From the sign of the eigenvalues and the direction of the vector field on the boundary of the phase space, we can lead to the conclusion that $D_{1},~{}D_{2}$ are unstable critical points, $D_{3},~{}D_{4}$ give saddle type nature and $D_{0}$ is the only global stable critical point. Here, $D_{0},~{}D_{3},~{}D_{4}$ correspond to accelerating points and $D_{1}$ belong to deceleration phase. As a result $D_{1}$ is the past matter dominated unstable critical point and $D_{0}$ is the late time completely DE dominated late time stable critical point. The phase space orbits are exhibited in the right plot of Fig. 8, providing the transition from the DM dominated accelerating phase to the DE dominated super accelerating phase analogous to the case $w<-1$ of the Model III.1.

### v2

@@ III.4 Model: $\Gamma=\Gamma_{0}H^{-1}$
In this series of matter creation models, one of the interesting matter creation rate is $\Gamma=\Gamma_{0}H^{-1}$, where $\Gamma_{0}$ is constant having dimension equal to the dimension of the square of the Hubble rate. In the following we describe the influence of this model on the two-fluid systems.
@@
where $\gamma=\Gamma_{0}/H_{0}^{2}$ is a dimensionless parameter and it is positive. In order to remove the singularity at $z=1$ of the above dynamical system, we introduce a new time variable $\tau$, defined through $dN=(1-z)^{2}d\tau$. The resulting autonomous system is topologically equivalent to (27) and is given by
$\frac{dx}{d\tau}=$ $x(1-x)\left(3w(1-z)^{2}+\gamma z^{2}\right),$ (28a)
$\frac{dz}{d\tau}=$ $\frac{3}{2}z(1-z)\left[\left(1+w(1-x)\right)(1-z)^{2}-\frac{\gamma xz^{2}}{3}\right].$ (28b)
If we look into the system (28), we find that x=0,x=1,z=0,z=1x=0,~x=1,~z=0,~z=1 are invariant manifolds. So, the domain $\bf R$} is positively invariant. For this model we obtain the decelerating parameter as $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\gamma xz^{2}}{(1-z)^{2}}\right]$. Thus, at present time (i.e. $z=1/2$) to obtain an accelerating phase of the universe the viable orbits should maintain the relation $(1-\gamma)x+(1+3w)(1-x)<0$. The critical points and different cosmological parameter at these points are depicted in the Table 4. Now we shall analyze the qualitative features of the system (28) for different values of $w$.
Critical point Existence Eigenvalue Stability $\mathbf{\Omega_{f}}$ $\mathbf{\Omega_{\rm dm}^{\rm eff}}$ $\mathbf{q}$ Acceleration
$D_{0}(0,0)$ Always $\left(3w,\frac{3}{2}(1+w)\right)$ Stable if w<1w<-1; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if 1<w<0-1<w<0;
@@
$D_{2}(0,1)$ Always $(\gamma,0)$ Non-hyperbolic Saddle if w>1w>-1; 1 0 Undefined Undetermined
Unstable if w<1w<-1
$D_{3}(1,1)$ Always $\left(-\gamma,\frac{\gamma}{2}\right)$ Always Saddle 0 1 -$\infty$ Yes
$D_{4}\left(1,\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$ $\gamma>0$ $\left(-\frac{\gamma(1+w)}{\left(1+\sqrt{\frac{\gamma}{3}}\right)^{2}},-\frac{\gamma}{\left(1+\sqrt{\frac{\gamma}{3}}\right)^{2}}\right)$ Stable if w>1w>-1 0 1 1-1 Yes
Saddle if w<1w<-1
$D_{5}\left(0,z_{c}\right)$ w=1w=-1 $\left(\gamma z_{c}^{2}-3(1-z_{c})^{2},0\right)$ Stable if $z_{c}<\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ 11 0 1-1 Yes
@@
$D_{6}\left(x_{c},\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$ w=1w=-1 $\left(0,-\frac{\gamma x_{c}}{\left(1+\sqrt{\frac{\gamma}{3}}\right)^{2}}\right)$ Stable $1-x_{c}$ $x_{c}$ 1-1 Yes
@@
Table 4: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (28) with the matter creation rate $\Gamma=\Gamma_{0}H^{-1}$.
@@
Figure 10: Description of the phase space controlled by the matter creation rate $\Gamma=\Gamma_{0}H^{-1}$. Upper Left Plot: The phase plot of the system (28) when we have assumed $w=0.1$ and $\gamma=2$. For other values of $w>0$ and $\gamma>0$, we can also obtain similar phase space structure. Upper Middle Plot: The phase space of the system (28) when the EoS $w$ takes the value 0. Here we use $\gamma=2$ but any positive value of $\gamma$ gives same type of phase portrait. Upper Right Plot: The phase plot of the system (28) considering w=0.8w=-0.8 and $\gamma=3$. Also, we can get similar type of graphics for any positive value of $\gamma$ and negative value of $w$ in the interval (1,0)(-1,0). Lower Left Plot: The phase space of the system (28) when the EoS $w$ takes the value 1-1. Here we use $\gamma=2$ but any positive value of $\gamma$ gives similar type of phase portrait. Lower Right Plot: The phase plot of the system (28) when we assume w=1.5w=-1.5 and $\gamma=3$. For other values of w<1w<-1 and $\gamma>0$, we can also obtain similar graphics. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with 1<q<0-1<q<0 and the yellow region corresponds to the super accelerating phase (i.e. q<1q<-1). The black dotted curve (in the lower left graph this curve is not visible because of the presence of the critical line) separating the cyan and yellow regions corresponds to q=1q=-1.
Figure 11: We display the evolution plots of the DM density parameter $\Omega_{\rm dm}^{\rm eff}$, the second fluid density parameter $\Omega_{f}$, the deceleration parameter $q$, and the effective equation of state of DM $w_{\rm dm}^{\rm eff}$ corresponding to the dynamical system (28), with the matter creation rate $\Gamma=\Gamma_{0}H^{-1}$. Upper Left Plot: This figure is drawn with the model parameters $w=0.1$, $\gamma=1$, and it shows that the final fate of the universe is late time accelerated evolution in cosmological constant era dominated by DM. Upper Right Plot: This figure is produced for the parameters values $w=0$ and $\gamma=1$. Here also, the final fate of the universe is late time accelerated evolution in cosmological constant era dominated by DM. Lower Left Plot: We have used w=1w=-1 and $\gamma=1$. This plot depicts that the ultimate fate of the universe is found to be late time accelerated evolution in cosmological constant epoch which is achieved after the occurrence of a decelerated DM dominated phase. Lower Right Plot: We have taken w=1.2w=-1.2 and $\gamma=1$. This plot highlights that the ultimate fate of the universe is found to be late time DE dominated accelerated evolution in phantom epoch which is achieved after the occurrence of a decelerated DM dominated phase.
[…]
@@
When $w>0$, we obtain five isolated critical points, namely, $D_{0},~D_{1},~D_{2},~D_{3}$ and $D_{4}$ in which $D_{0},~D_{2}$ are dominated by the second fluid only ($\Omega_{f}=1$) and $D_{1},~D_{3},~D_{4}$ are DM dominated points ($\Omega_{\rm dm}=1$). Inspecting the nature of eigenvalues from Table 4, according to the linear stability analysis, we find that $D_{0}$ is unstable, $D_{1},~D_{3}$ both are saddle by nature and $D_{4}$ is the only globally stable point. The critical point $D_{2}$ at which decelerating parameter is undefined, behaves like saddle point because along $z$ axis, $z$ is increasing and along $z=1$ line $x$ is increasing. Here, the points $D_{0},~D_{1}$ lie in the decelerating phase and $D_{3},D_{4}$ also belong to the accelerated phase of the universe. The phase space trajectories are highlighted in the upper left plot of the Fig. 10, indicating a transition of our universe from the second fluid dominated decelerating phase to a DM dominated accelerating phase where the slowing down of the cosmic acceleration is observed similar to what we have already seen for the two-fluid system of the Model $\Gamma=\Gamma_{0}$ with $w>0$, see section III.1. This is further supported by the upper left plot of Fig. 11, which illustrates the evolution of $\Omega_{\rm dm}^{\rm eff}$, $\Omega_{f}$, $q$ and $w_{\rm dm}^{\rm eff}$.
@@
When $w=0$, in this case, we obtain total three isolated critical points $D_{2},~D_{3},~D_{4}$ and one critical line $D_{7}(x_{c},0)$. In our phase space, $x^{\prime}$ is always positive and, $z$ is increasing for $z<\frac{1}{1+\sqrt{\frac{\gamma}{3}}\sqrt{x}}$ and it is decreasing for $z>\frac{1}{1+\sqrt{\frac{\gamma}{3}}\sqrt{x}}$. Therefore, $D_{7}$ where DM and DE coexist, is always unstable and $D_{2},~D_{3}$ give saddle type behavior. Since in our domain $D_{4}$ is the only stable point, we claim that it is a globally stable point. The accelerated phase includes $D_{3},~D_{4}$ and on the other side, the decelerated phase comprises the line $D_{7}$. The phase space orbits are depicted in the upper middle plot of the Fig. 10, showing an alteration of the universe from the decelerating scaling solutions (DM and second fluid both exist in the picture) to DM dominated accelerating era where the slowing down of the cosmic acceleration is noted - the same feature has been observed in the model $\Gamma=\Gamma_{0}$ with $w=0$, see section III.1. As supporting evidence, the upper right plot in Fig. 11 shows the evolution of the density parameters $\Omega_{\rm dm}^{\rm eff}$, $\Omega_{f}$, $q$ and $w_{\rm dm}^{\rm eff}$.
@@
When $\frac{1}{3}\leq w<0$, once again we can get five isolated critical points, namely, $D_{0},~D_{1},~D_{2},~D_{3}$ and $D_{4}$ by solving the autonomous system (28). Looking into the eigenvalues from the Table 3 and using the direction of the flow on the boundary of phase space, one can easily conclude about the stability of the critical points. Therefore, $D_{0},~D_{2}$ and $D_{3}$ give saddle type character, $D_{1}$ shows unstable nature and as before $D_{4}$ corresponds to matter dominated late time globally stable point. Here also, $D_{0}$, $D_{1}$ lie in the decelerating phase and $D_{3}$ belongs to the accelerating phase. The evolution of the solution curves are clearly observed in the upper right plot of the Fig. 10, highlighting the alteration from DM dominated decelerating epoch to DM dominated accelerating epoch with its slowing down nature, which one can meanwhile find in the Model III.1 with the case 1/3<w<0-1/3<w<0. The case with $1<w<-\frac{1}{3}$ properly replicates the above case where the range of $w$ is $\frac{1}{3}\leq w<0$. The only difference lies on the decelerating parameter at the point $D_{0}$ which shows acceleration. Thus, the phase plot is the upper right plot of Fig. 10.
@@
For w=1w=-1, we have a set of critical points which contains two isolated critical points $D_{1},~D_{3}$ and two critical lines $D_{5}(0,z_{c})$, $D_{6}\left(x_{c},\frac{1}{1+\sqrt{\frac{\gamma}{3}}}\right)$. In the phase space above the line $z=\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$, $x$ and $z$ show increasing and decreasing character, respectively, and below this line, $x$ is decreasing function while $z$ is increasing function. The equation of the separatrix joining the points $D_{1}$ and $\left(0,\frac{1}{\sqrt{\frac{\gamma}{3}}}\right)$ is $z=\sqrt{\frac{3}{\gamma}}(1-z)\sqrt{1-x}$ which separates the phase space below the line $z=\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ into two parts; one is below the separatrix and other one is the above the separatrix. In the first part, any trajectory approaches to the line $D_{5}$ and in the second part, trajectories converge to the line $D_{6}$. Therefore, $D_{6}$ where DM and DE coexist (except the end points), gives stable like nature. Here, $D_{1}$ and $D_{6}$ with $z>\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ show unstable nature and $D_{3}$ is a saddle point. The accelerating phase includes $D_{3}$, $D_{5}$, $D_{6}$ and $D_{1}$ corresponds to deceleration. The part of the line $D_{5}$ below $z=\frac{1}{1+\sqrt{\frac{\gamma}{3}}}$ corresponds to late time completely DE dominated stable points. On the other hand, the line $D_{6}$ has the ability to solve the coincidence problem. The phase space structure is properly presented in the lower left plot of Fig. 10 which depicts two types of transitions: the first one is DM dominated decelerating phase to accelerating scaling solutions (DM and DE both exist) and the second one is DM dominated decelerating phase to completely DE dominated accelerating phase, that is similar to the case w=1w=-1 of Model III.1. The lower left plot in Fig. 11 supports this scenario by displaying the evolution of $\Omega_{\rm dm}^{\rm eff}$, $\Omega_{f}$, $q$ and $w_{\rm dm}^{\rm eff}$.
@@
When w<1w<-1, we again find five critical points, namely, $D_{0},~D_{1},~D_{2},~D_{3},~D_{4}$ which lie on the boundary of the phase space region $R$. From the sign of the eigenvalues and the direction of the vector field on the boundary of the phase space, we can lead to the conclusion that $D_{1},~D_{2}$ are unstable critical points, $D_{3},~D_{4}$ give saddle type nature and $D_{0}$ is the only global stable critical point. Here, $D_{0},~D_{3},~D_{4}$ correspond to accelerating points and $D_{1}$ belong to deceleration phase. As a result $D_{1}$ is the past matter dominated unstable critical point and $D_{0}$ is the late time completely DE dominated late time stable critical point. The phase space orbits are exhibited in the lower right plot of Fig. 10, providing the transition from the DM dominated accelerating phase to the DE dominated super accelerating phase analogous to the case w<1w<-1 of the Model III.1. As shown in the lower right plot of Fig. 11, the evolution of $\Omega_{\rm dm}^{\rm eff}$, $\Omega_{f}$, $q$, and $w_{\rm dm}^{\rm eff}$ reinforces this behavior.
Finally, in Fig. 12, we illustrate the evolution of $H/H_{0}$ for the matter creation scenario with $\Gamma=\Gamma_{0}H^{-1}$ compared to $\Lambda{\rm CDM}$. The fractional difference $\Delta H/H$ is shown in the lower panel. We find that the models match at low redshift, while at higher-redshift, differences highlight the role of particle creation.